at full speed, a motor boat can travel 32 miles upstream in 8 hours, but it can travle the same distance downstram in 4 hours. What is the rate of the current in miles per hour.
Answer by stanbon(57282)

At full speed, a motor boat can travel 32 miles upstream in 8 hours, but it can travel the same distance downstram in 4 hours. What is the rate of the current in miles per hr.
---------------------------
Upstream DATA:
distance = 32 miles; time = 8 hrs ; rate = 32/8 = 4 mph
-----
Downstream DATA:
distance = 32 miles ; time = 4 hrs ; rate = 32/4 8 mph
-----
Equation:
Let boat speed = b
Let current speed = c
---
b-c = 4
b+c = 8
---
2b = 12
b = 6 mph (speed of the boat in still water)\
----
Since b + c = 8, c = 2 mph
=============================
cheers,
Stan H.
=============================